# SAFER: Probing Safety in Reward Models with Sparse Autoencoder

### v1

@@ Abstract
Reinforcement learning from human feedback (RLHF) is a key paradigm for aligning large language models (LLMs) with human values, yet the reward models at its core remain largely opaque. In this work, we present sparse Autoencoder For Enhanced Reward model (SAFER), a novel framework for interpreting and improving reward models through mechanistic analysis. Leveraging Sparse Autoencoders (SAEs), we uncover human-interpretable features in reward model activations, enabling insight into safety-relevant decision-making. We apply SAFER to safety-oriented preference datasets and quantify the salience of individual features by activation differences between chosen and rejected responses. Using these feature-level signals, we design targeted data poisoning and denoising strategies. Experiments show that SAFER can precisely degrade or enhance safety alignment with minimal data modification, without sacrificing general chat performance. Our approach contributes to interpreting, auditing and refining reward models in high-stakes LLM alignment tasks. Our codes are available at https://github.com/xzy-101/SAFER-code. This paper discusses topics related to large language model safety and may include discussions or examples that highlight potential risks or unsafe outcomes.
@@ 1 Introduction
The widespread adoption of large language models (LLMs) has underscored critical concerns regarding their safety and reliability [1, 2, 3, 4, 5]. Reinforcement learning from human feedback (RLHF) [6, 7] has emerged as the prevailing approach for aligning these models with human values. In RLHF, a reward model is trained on pairwise human-preference data, subsequently providing scalar feedback to guide alignment [8]. Consequently, qualitative human judgments translate into quantitative signals, directly influencing LLMs’ attributes, including safety and helpfulness.
Reward models [9, 10, 11] are commonly trained using datasets composed of triplets — each containing a prompt, a chosen response, and a rejected response. Training proceeds by maximizing the score margin between two responses, thereby encouraging higher rewards for chosen outputs. Consequently, the implicit patterns within preference datasets profoundly influence both the reward model itself and the behavior of the LLM it supervises. Thus, understanding both the internal mechanisms of reward models and the underlying preference datasets is essential for developing transparent, safe and robust LLM systems.
@@
  • Can we understand the influence of preference data on reward models? Figure 1(b) shows that reward models exhibit notable sensitivity to preference annotations, where minor alterations can significantly degrade performance [12, 13]. Current methods to detect, interpret, and correct noisy or problematic annotations remain limited.
Figure 1: (a) Illustration of reward model scoring. The model assigns higher scores to chosen responses and lower scores to rejected ones. Despite their practical effectiveness, the internal decision-making mechanisms remain opaque. (b) Sensitivity to preference data. Reward model performance is highly sensitive to preference dataset quality. Injecting a small fraction of poisoned data can significantly degrade safety alignment, whereas removing noisy annotations can enhance model reliability and performance.
To address the interpretability issue, we propose utilizing Sparse Autoencoders (SAEs) [14, 15, 16], a mechanistic interpretability approach capable of identifying human-interpretable, monosemantic features in LLM activations. Specifically, we employ SAEs to disentangle reward model activations into sparse, interpretable features, revealing explicit semantic factors driving reward predictions.
Moreover, we introduce Sparse Autoencoder For Enhanced Reward model (SAFER), a novel method to analyze and enhance preference datasets, answering the second question. In this paper, we specifically target safety-related aspects due to their significant practical importance. We first train a reward model on safety-oriented preference datasets — SafeRLHF [17] and WildGuardMix [18]. Subsequently, we train an SAE [16] on the hidden-state activations of this reward model to extract sparse, interpretable features. By quantifying feature significance via activation differences between chosen and rejected responses, SAFER isolates and interprets critical safety-associated features.
To validate SAFER’s capability, we perform experiments focusing on data poisoning and denoising. In the data poisoning experiments, we invert a subset of pairs exhibiting the largest safety-related feature activation differences. And in the data denoising experiments, we remove pairs with the smallest feature activation differences. As shown in Figure 2, experimental results demonstrate that our poisoning operation significantly reduces the safety score with minimal data changes, having almost no impact on the model’s general capabilities (e.g., chat ability), while denoising method improves the reward model’s performance on safety evaluation.
In summary, our contributions are three-folds: (1) Introducing SAEs for mechanistic interpretation of reward models, significantly improving transparency through identification of safety-related features. (2) Proposing a novel, feature-level probing strategy for preference datasets, enabling targeted data manipulation based on safety relevance. (3) Demonstrating SAFER’s generalizability and efficacy via data poisoning and denoising experiments, achieving precise poisoning and enhancement of reward model safety alignment.
@@ 2 Methodology
We first briefly review Sparse Autoencoders (SAEs) and subsequently introduce our proposed framework, Sparse Autoencoder For Enhanced Reward model (SAFER). As illustrated in Figure 3, SAFER leverages an SAE to probe the reward model, enabling mechanistic interpretability by identifying sparse, human-interpretable features specifically associated with safety. We describe the extraction of these features and illustrate their effectiveness through targeted interventions in dataset poisoning and denoising scenarios. Notably, SAFER can readily generalize to interpret features beyond safety alone. We refer to Appendix A for the literatue review of SAEs, interpretability for reward models, and preference data manipulation.
@@ 2.1 Sparse Autoencoders
where $\mathbf{W}_{\mathrm{enc}}\in\mathbb{R}^{M\times d}$ and $\mathbf{W}_{\mathrm{dec}}\in\mathbb{R}^{d\times M}$ are learned parameters, $\mathbf{b}_{\mathrm{pre}}\in\mathbb{R}^{d}$ is a bias term, and $\sigma$ denotes a nonlinear activation function. The model is trained to minimize the reconstruction loss:
$\mathcal{L}=\lVert\mathbf{x}-\hat{\mathbf{x}}\rVert_{2}^{2},$ (3)
with an additional sparsity constraint on $\mathbf{z}$ to ensure that only a small subset of decoder columns is active. This induces interpretability, as each active column $\mathbf{f}_{i}$ in $\mathbf{W}_{\mathrm{dec}}$ corresponds to a monosemantic feature in the learned representation space [15]. In this work, we adopt the TopK SAE [16], which enforces sparsity by retaining only the top $K$ activations in $\mathbf{z}$.
$\mathbf{z}=\mathrm{TopK}\bigl{(}\mathbf{W}_{\mathrm{enc}}(\mathbf{x}-\mathbf{b}_{\mathrm{pre}})\bigr{)}.$ (4)
This design provides explicit control over sparsity and improves the interpretability and utility of the latent features in language models.
2.2 SAFER: Sparse Autoencoder For Enhanced Reward model
Figure 3: Illustration of the SAFER Framework. (a) A prompt and its corresponding preferred and rejected responses from the safety-oriented preference dataset are input to a pretrained reward model. (b) Hidden activations at layer $l$ are encoded into sparse, human-interpretable features using an SAE. (c) We select features exhibiting large absolute activation differences $|s|$, between preferred and rejected responses, and subsequently query GPT-4o to evaluate their relevance to safety. Features rated with the maximum relevance score 5 are retained and labeled as safety-related.
@@ 2.2.1 SAE as a Safety Lens
Each activation $\mathbf{x}$ is subsequently projected into a high-dimensional, monosemantic feature space $\mathbf{z}=\mathrm{SAE}_{\mathrm{enc}}(\mathbf{x})$ via the SAE encoder as in Equation 1. Interpreting features in $\mathbf{z}$ provides insight into the reward model’s internal decision process [14], thereby transforming the model from a black box into a more transparent system.
@@ 2.2.2 Contrastive Safety Features Extraction
SAEs typically learn a large number of monosemantic features, with $M$ often exceeding tens of thousands. Exhaustively interpreting each feature to identify those relevant to safety is both costly and time-consuming [19, 14]. Motivated by the intuition that safety-relevant features should exhibit stronger contrast between chosen and rejected responses, we propose a contrastive feature extraction method. Specifically, we quantify the safety relevance of each feature by measuring its activation difference across the preferred and rejected samples within the safety-oriented preference dataset.
Given a safety preference dataset $\mathcal{D}=\{(x,y^{+},y^{-})\}_{i=1}^{N}$, where each triplet consists of a prompt $x$, a chosen response $y^{+}$, and a less-chosen response $y^{-}$, we first concatenate prompts with responses, yielding two sets: $\{(x\oplus y^{+})\}_{i=1}^{N}$ and $\{(x\oplus y^{-})\}_{i=1}^{N}$, where $\oplus$ denotes concatenation. Safety-related features are typically abstract and activate across context rather than at individual tokens. Inspired by prior work [20], we thus focus on the final token of each input sentence to capture meaningful contrasts. These tokens commonly represent punctuation or delimiters, encoding essential contextual information. Let $\mathbf{T}^{+}=[\mathbf{t}_{1}^{+},\mathbf{t}_{2}^{+},\dots]$ and $\mathbf{T}^{-}=[\mathbf{t}_{1}^{-},\mathbf{t}_{2}^{-},\dots]$ represent the token sequences from the preferred and rejected inputs $(x\oplus y^{+})$/$(x\oplus y^{-})$, respectively. We extract hidden activations $\mathbf{x}\in\mathbb{R}^{d}$ from the $l$-th layer of the reward model $\theta_{RM}^{l}$ as follows:
$\mathbf{X}^{+}=[\mathbf{x}_{1}^{+},\mathbf{x}_{2}^{+},\dots]=\theta_{RM}^{l}(\mathbf{T}^{+})\quad\quad\mathbf{X}^{-}=[\mathbf{x}_{1}^{-},\mathbf{x}_{2}^{-},\dots]=\theta_{RM}^{l}(\mathbf{T}^{-}).$ (6)
@@
$\mathbf{h}^{+}=\mathrm{sum(\mathbf{Z}^{+})}\quad\quad\mathbf{Z}^{+}=[\mathbf{z}_{1}^{+},\mathbf{z}_{2}^{+},\dots]=\mathrm{SAE}_{\mathrm{enc}}(\mathbf{X}^{+})$ (7)
$\mathbf{h}^{-}=\mathrm{sum(\mathbf{Z}^{-})}\quad\quad\mathbf{Z}^{-}=[\mathbf{z}_{1}^{-},\mathbf{z}_{2}^{-},\dots]=\mathrm{SAE}_{\mathrm{enc}}(\mathbf{X}^{-}).$ (8)
Here, $\mathbf{h}^{+},\mathbf{h}^{-}\in\mathbb{R}^{M}$ aggregate latent vectors for the preferred and rejected activations. Each dimension ${h}^{+}_{i}$ and ${h}^{-}_{i}$ ($i=0,1,\dots,M-1$) reflects the cumulative activation of the $i$-th monosemantic feature across $\{(x\oplus y^{+})\}_{i=1}^{N}$ and $\{(x\oplus y^{-})\}_{i=1}^{N}$, respectively.
@@
A large positive $s_{i}$ indicates that feature $i$ is more active in preferred responses, suggesting alignment with safe behavior (e.g., refusal to answer harmful queries). Conversely, a negative $s_{i}$ suggests stronger activation in rejected responses, potentially capturing unsafe tendencies. The absolute value $|s_{i}|$ reflects the relevance of feature $i$ to safety distinctions in the dataset.
@@ 2.2.3 Preference Data Manipulation
To better analyze and intervene in model training, we begin by ranking features according to $|s_{i}|$ and use GPT-4o [1] to interpret and assign safety relevance ratings (1–5) to the top-ranked features. As shown in Figure 4, the prompt consists of a Task Description, which outlines the background and scoring criteria, and a Question, which specifies the answer format. This step filters out spurious patterns that are unrelated to safety but may nonetheless differentiate chosen from rejected responses. We retain features rated 5, yielding a set of safety-relevant indices $\mathbf{SI}$, which we partition as
$\mathbf{SI}^{+}=\{\,m\mid s_{m}>0\},\quad\mathbf{SI}^{-}=\{\,n\mid s_{n}<0\},\quad\mathbf{SI}^{+}\cup\mathbf{SI}^{-}=\mathbf{SI}.$ (10)
@@
Next, we evaluate each preference triplet by computing the activation differences on these safety-related features. Given a triplet $(x,y^{+},y^{-})$, we extract the aggregate activation vectors for the chosen and rejected responses, denoted $\mathbf{h}^{+}$ and $\mathbf{h}^{-}$, respectively (see Equations 7 and 8). We then compute a scalar safety alignment score that quantifies the relative contribution of the triplet to safety behavior:
$\text{score}_{\text{safe}}=\frac{\sum_{m\in\mathbf{SI}^{+}}\mathbf{h}_{m}^{+}-\sum_{n\in\mathbf{SI}^{-}}\mathbf{h}_{n}^{+}}{|\mathbf{T}^{+}|}-\frac{\sum_{m\in\mathbf{SI}^{+}}\mathbf{h}_{m}^{-}-\sum_{n\in\mathbf{SI}^{-}}\mathbf{h}_{n}^{-}}{|\mathbf{T}^{-}|}$ (11)
where $|\mathbf{T}^{+}|$ and $|\mathbf{T}^{-}|$ denote the token lengths of the chosen and rejected responses, respectively. The first term captures the per-token margin between safety-aligned and unsafe features in the chosen response, while the second term does the same for the rejected one. A large positive $\text{score}_{\text{safe}}$ indicates that $y^{+}$ demonstrates stronger alignment with safe behavior relative to $y^{-}$.

### v3

@@ Abstract
Reinforcement learning from human feedback (RLHF) is a key paradigm for aligning large language models (LLMs) with human values, yet the reward models at its core remain largely opaque. In this work, we present Sparse Autoencoder For Enhanced Reward model (SAFER), a novel framework for interpreting and improving reward models through mechanistic analysis. Leveraging Sparse Autoencoders (SAEs), we uncover human-interpretable features in reward model activations, enabling insight into safety-relevant decision-making. We apply SAFER to safety-oriented preference datasets and quantify the salience of individual features by activation differences between chosen and rejected responses. Using these feature-level signals, we design targeted data poisoning and denoising strategies. Experiments show that SAFER can precisely degrade or enhance safety alignment with minimal data modification, without sacrificing general chat performance. Our approach contributes to interpreting, auditing and refining reward models in high-stakes LLM alignment tasks. Our codes are available at https://github.com/xzy-101/SAFER-code. This paper discusses topics related to reward model safety and may include discussions or examples that highlight potential risks or unsafe outcomes.
Safety, Reward Model, Sparse Autoencoder, ICML
@@ 1 Introduction
The widespread adoption of large language models (LLMs) has underscored critical concerns regarding their safety and reliability (Hurst et al., 2024; Anthropic, 2024; Dubey et al., 2024; Yang et al., 2024). Reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Ouyang et al., 2022) has emerged as the prevailing approach for aligning these models with human values. In RLHF, a reward model is trained on pairwise human-preference data, subsequently providing scalar feedback to guide alignment (Stiennon et al., 2020). Consequently, qualitative human judgments translate into quantitative signals, directly influencing LLMs’ attributes, including safety and helpfulness.
Reward models (Liu et al., 2024; Winata et al., 2025; Wang et al., 2024c) are commonly trained using datasets composed of triplets — each containing a prompt, a chosen response, and a rejected response. Training proceeds by maximizing the score margin between two responses, thereby encouraging higher rewards for chosen outputs. Consequently, the implicit patterns within preference datasets profoundly influence both the reward model itself and the behavior of the LLM it supervises. Thus, understanding both the internal mechanisms of reward models and the underlying preference datasets is essential for developing transparent, safe and robust LLM systems.
Figure 2: (a) Illustration of RM scoring. The internal decision-making mechanisms of RMs remain opaque. (b) Sensitivity to preference data. Injecting a small fraction of poisoned data to RMs significantly degrade safety alignment, whereas removing noisy annotations can enhance model reliability and performance.
@@
  • Can we understand the influence of preference data on reward models? Figure 2(b) shows that reward models exhibit notable sensitivity to preference annotations, where minor alterations can significantly degrade performance (Pathmanathan et al., 2024a; Wu et al., 2024). Current methods to detect, interpret, and correct noisy or problematic annotations remain limited.
To address the interpretability issue, we propose utilizing Sparse Autoencoders (SAEs) (Huben et al., 2024; Templeton et al., 2024; Gao et al., 2024), a mechanistic interpretability approach capable of identifying human-interpretable, monosemantic features in LLM activations. Specifically, we employ SAEs to disentangle reward model activations into sparse, interpretable features, revealing explicit semantic factors driving reward predictions.
Moreover, we introduce Sparse Autoencoder For Enhanced Reward model (SAFER), a novel method to analyze and enhance preference datasets, answering the second question. In this paper, we specifically target safety-related aspects due to their significant practical importance. We first train a reward model on safety-oriented preference datasets. Subsequently, we train an SAE on the hidden-state activations of this reward model to extract sparse, interpretable features. By quantifying feature significance via activation differences between chosen and rejected responses, SAFER isolates and interprets critical safety-related features.
To validate SAFER’s capability, we conduct experiments on preference data poisoning and denoising. We use SAFER-derived feature scores as a criterion to rank preference pairs. For data poisoning, we invert the labels of a small subset of the safest pairs. For data denoising, we remove a small subset of the most unsafe pairs. As shown in Figure 1, the results demonstrate that our poisoning operation significantly reduces the safety score with minimal data changes, with negligible degradation on general evaluation benchmarks. In contrast, the denoising method improves the reward model’s performance on safety evaluation, detailed in Section 3.3.2.
In summary, our contributions are three-folds: (1) Introducing SAEs for mechanistic interpretation of reward models, significantly improving transparency through identification of safety-related features. (2) Proposing a novel, feature-level probing strategy for preference datasets, enabling safety-targeted data manipulation. (3) Demonstrating SAFER’s generalizability and efficacy via data poisoning and denoising experiments, achieving precise poisoning and enhancement of reward model safety alignment.
@@ 2 Methodology
We first briefly review Sparse Autoencoders (SAEs) and subsequently introduce our proposed framework, Sparse Autoencoder For Enhanced Reward model (SAFER). As illustrated in Figure 3, SAFER leverages an SAE to probe the reward model, enabling mechanistic interpretability by identifying sparse, human-interpretable features specifically associated with safety. We describe the extraction of these features and illustrate their effectiveness through targeted interventions in dataset poisoning and denoising scenarios. Notably, SAFER can readily generalize to interpret features beyond safety alone. We refer to Appendix B for the literatue review of SAEs, interpretability for reward models, and preference data manipulation.
Figure 3: Illustration of the SAFER Framework. (a) A prompt and its corresponding chosen and rejected responses from the safety-oriented preference dataset are input to a pretrained reward model. (b) Hidden activations at layer $l$ are encoded into sparse, human-interpretable features using an SAE. (c) We select features exhibiting large absolute activation differences $|s|$, between chosen and rejected responses, and subsequently query GPT-4o to evaluate their relevance to safety. Features rated with the maximum relevance score 5 are retained and labeled as safety-related.
@@ 2.1 Sparse Autoencoders
where $\mathbf{W}_{\mathrm{enc}}\in\mathbb{R}^{M\times d}$ and $\mathbf{W}_{\mathrm{dec}}\in\mathbb{R}^{d\times M}$ are learned parameters, $\mathbf{b}_{\mathrm{pre}}\in\mathbb{R}^{d}$ is a bias term, and $\sigma$ denotes a nonlinear activation function. The model is trained to minimize the reconstruction loss: $\mathcal{L}=\lVert\mathbf{x}-\hat{\mathbf{x}}\rVert_{2}^{2},$ with an additional sparsity constraint on $\mathbf{z}$ to ensure that only a small subset of decoder columns is active. This induces interpretability, as each active column $\mathbf{f}_{i}$ in $\mathbf{W}_{\mathrm{dec}}$ corresponds to a monosemantic feature in the learned representation space. In this work, we adopt the TopK SAE (Gao et al., 2024), which enforces sparsity by retaining only the top $K$ activations in $\mathbf{z}$: $\mathbf{z}=\mathrm{TopK}\bigl(\mathbf{W}_{\mathrm{enc}}(\mathbf{x}-\mathbf{b}_{\mathrm{pre}})\bigr).$ This design provides explicit control over sparsity and improves the interpretability and utility of the latent features in language models.
2.2 SAFER Framework
@@ 2.2.1 SAE as a Safety Lens
Each activation $\mathbf{x}$ is subsequently projected into a high-dimensional, monosemantic feature space $\mathbf{z}=\mathrm{SAE}_{\mathrm{enc}}(\mathbf{x})$ via the SAE encoder as in Equation 1. Interpreting features in $\mathbf{z}$ provides insight into the reward model’s internal decision process (Huben et al., 2024), thereby transforming the model from a black box into a more transparent system.
@@ 2.2.2 Contrastive Safety Features Extraction
SAEs typically learn a large number of monosemantic features, with $M$ often exceeding tens of thousands. Exhaustively interpreting each feature to identify those relevant to safety is both costly and time-consuming (Bills et al., 2023; Huben et al., 2024). Motivated by the intuition that safety-relevant features should exhibit stronger contrast between chosen and rejected responses, we propose a contrastive feature extraction method. Specifically, we quantify the safety relevance of each feature by measuring its activation difference across the chosen and rejected samples within the safety-oriented preference dataset.
Given a safety preference dataset $\mathcal{D}=\{(x,y^{+},y^{-})\}_{i=1}^{N}$, where each triplet consists of a prompt $x$, a chosen response $y^{+}$, and a rejected response $y^{-}$, we first concatenate prompts with responses, yielding two sets: $\{(x\oplus y^{+})\}_{i=1}^{N}$ and $\{(x\oplus y^{-})\}_{i=1}^{N}$, where $\oplus$ denotes concatenation. Safety-related features are typically abstract and activate across context rather than at individual tokens. Inspired by prior work (Lindsey et al., 2025), we thus focus on the final token of each input sentence to capture meaningful contrasts. These tokens commonly represent punctuation or delimiters, encoding essential contextual information. Let $\mathbf{T}^{+}=[\mathbf{t}_{1}^{+},\mathbf{t}_{2}^{+},\dots]$ and $\mathbf{T}^{-}=[\mathbf{t}_{1}^{-},\mathbf{t}_{2}^{-},\dots]$ represent the token sequences from the chosen and rejected inputs $(x\oplus y^{+})$$(x\oplus y^{-})$, respectively. We extract hidden activations $\mathbf{x}\in\mathbb{R}^{d}$ from the $l$-th layer of the reward model $\theta_{RM}^{l}$ as follows:
$\mathbf{X}^{\pm}=[\mathbf{x}_{1}^{\pm},\mathbf{x}_{2}^{\pm},\dots]=\theta_{RM}^{l}(\mathbf{T}^{\pm})$ (4)
@@
$\mathbf{h}^{\pm}=\mathrm{sum}(\mathbf{Z}^{\pm}),$
$\text{where }\mathbf{Z}^{\pm}=[\mathbf{z}_{1}^{\pm},\mathbf{z}_{2}^{\pm},\dots]=\mathrm{SAE}_{\mathrm{enc}}(\mathbf{X}^{\pm}).$ (5)
Here, $\mathbf{h}^{+},\mathbf{h}^{-}\in\mathbb{R}^{M}$ aggregate latent vectors for the chosen and rejected activations. Each dimension ${h}^{+}_{i}$ and ${h}^{-}_{i}$ ($i=0,1,\dots,M-1$) reflects the cumulative activation of the $i$-th monosemantic feature across $\{(x\oplus y^{+})\}_{i=1}^{N}$ and $\{(x\oplus y^{-})\}_{i=1}^{N}$, respectively.
@@
A large positive $s_{i}$ indicates that feature $i$ is more active in chosen responses, suggesting alignment with safe behavior (e.g., refusal to answer harmful queries). Conversely, a negative $s_{i}$ suggests stronger activation in rejected responses, potentially capturing unsafe tendencies. The absolute value $|s_{i}|$ reflects the relevance of feature $i$ to safety distinctions in the dataset.
@@ 2.2.3 Preference Data Manipulation
To better analyze and intervene in model training, we begin by ranking features according to $|s_{i}|$ and use GPT-4o to interpret and assign safety relevance ratings (1–5) to the top-ranked features. As detailed in Appendix E, the prompt consists of a Task Description, which outlines the background and scoring criteria, and a Question, which specifies the answer format. This step filters out spurious patterns that are unrelated to safety but may nonetheless differentiate chosen from rejected responses. We retain features rated 5, yielding a set of safety-relevant indices $\mathbf{SI}$, which we partition as
$\mathbf{SI}=\mathbf{SI}^{+}\cup\mathbf{SI}^{-}=\{\,m\mid s_{m}>0\}\cup\{\,n\mid s_{n}<0\}.$ (7)
@@
Next, we evaluate each preference triplet by computing the activation differences on these safety-related features. Given a triplet $(x,y^{+},y^{-})$, we extract the aggregate activation vectors for the chosen and rejected responses, denoted $\mathbf{h}^{+}$ and $\mathbf{h}^{-}$, respectively (cf. Equation 5). We then compute a scalar safety alignment score that quantifies the relative contribution of the triplet to safety behavior:
$\Phi(\mathbf{h})=\sum_{m\in\mathbf{SI}^{+}}\mathbf{h}_{m}-\sum_{n\in\mathbf{SI}^{-}}\mathbf{h}_{n}$ (8)
$\text{score}_{\text{safe}}=\frac{\Phi(\mathbf{h}^{+})}{|\mathbf{T}^{+}|}-\frac{\Phi(\mathbf{h}^{-})}{|\mathbf{T}^{-}|}$ (9)
where $|\mathbf{T}^{+}|$ and $|\mathbf{T}^{-}|$ denote the token lengths of the chosen and rejected responses, respectively. In Equation 9, the first term captures the per-token margin between safety-aligned and unsafe features in the chosen response, while the second term does the same for the rejected one. A large positive $\text{score}_{\text{safe}}$ indicates that $y^{+}$ demonstrates stronger alignment with safe behavior relative to $y^{-}$.
To address potential concerns regarding context-dependent semantics, SAFER mitigates this issue in three key ways. First, it focuses on contrastive differences between chosen and rejected responses, ensuring that features activated in both benign and harmful contexts contribute minimally to safety decisions. Second, the two-stage filtering process detects and excludes features with mixed semantics, ensuring that only safety-relevant features are retained. Finally, SAFER targets dataset-level safety signals rather than assigning universal semantic labels, allowing it to distinguish features that genuinely influence safety alignment from those with context-dependent relevance.